OpenStudy (anonymous):

Calculate the surface area of the circle (πd2/4): 3.14*5^2/4=19.6 Is that right? If it is- Calculate the surface area of the dishwashing circle to molecules per layer (this requires converting the surface area from meters to nanometers):

OpenStudy (chmvijay):

what is ur d LOL:)

OpenStudy (anonymous):

5 cm

OpenStudy (chmvijay):

Equation is right ? then u r right

OpenStudy (anonymous):

YaY! I just found it hard to grasp that it went from 5 to 19.6

OpenStudy (chmvijay):

Area= Pie R^2 but R =d/2 area= Pie R^2/4 ur right LOL:)

OpenStudy (chmvijay):

5 is diameter or radius LOL:) if it is diameter ur right

OpenStudy (anonymous):

diameter lol thank you. can you help me with the second part?

OpenStudy (chmvijay):

its maths question LOL:) why are u posting it here ???

OpenStudy (anonymous):

because it was from a chemistry lab

OpenStudy (chmvijay):

LOL:)